2 buses leave at the same time from 2 terminals 792 km apart and meet in 9 hrs. one travels 8 kph faster than the othe. how far did the faster bus travels
Answer by checkley71(8403)

9S+9(S+8)=792
9S+9S+72=792
18S=792-72
18S=720
S=720/18
S=40 KPH FOR THE SLOWER BUS
40+8=48 KPH FOR THE FASTER BUS.
PROOF
9*40+9*48=792
360+432=792
792=792